OpenStudy (anonymous):

A sparrow flies to see a friend at a speed of 4 km/h. His friend is out, so the sparrow immediately returns home at a speed of 5km/h. The complete journey took 54 minutes. How far away does his friend live?

OpenStudy (tkhunny):

D = Rate * Time -- Every Time Go!

OpenStudy (gabebae):

let distance be x km while going to meet 4km/h time taken = x/4 return speed = 5km/h time = x/5 x/5 +x/4 = 54/60 (4x+5x)/20=54/60 9x=54*20/60 9x=18 /9 x= 2km Answer is 2km

OpenStudy (anonymous):

I know the answer but I need explanation...

OpenStudy (gabebae):

I just gave you it?

OpenStudy (tkhunny):

And thus we see that simply dumping your answer is an insufficient response. It's also contrary to the Code of Conduct.

OpenStudy (anonymous):

x/5 +x/4 = 54/60 how from this became (4x+5x)/20=54/60 this

OpenStudy (gabebae):

I explained it though

OpenStudy (anonymous):

GabeBae by the way your answer is copied from another site so i did not see smth new from you

OpenStudy (tkhunny):

Distance = Rate * Time -- This is what you need. Let's use Gabe's definition: x is the Distance <== Very important to WRITE THIS DOWN. Trip Out: x = Rate * Time = (4 km/h)*(Time Out) Notice the very clear and direct usage of the basic formula, Distance = Rate * Time Trip Back: x = Rate * Time = (5 km/h)*(Time Back) Notice the very clear and direct usage of the basic formula, Distance = Rate * Time We also know that (Time Out) + (Time Back) = 54 min = (54/60) hr All that remains is for you to demonstrate your best algebra skills.

OpenStudy (anonymous):

Thank you so much Mr. Tkhunny! That is the best professional explanation.

OpenStudy (tkhunny):

Was this a test of explanations? We like Gabe, too. People who are willing to help are very valuable. A little practice might improve the presentation, but we do NOT want to discourage the effort.

OpenStudy (gabebae):

Google is a very close friend of mine :)

OpenStudy (tkhunny):

@GabeBae That's good. It is a valuable resource. You should use it to develop your own style. :-)

OpenStudy (gabebae):

Style? I'm more of a science person but i'm not taking science for this semester